OpenStudy (anonymous):

Solve the following equation. 4(3 - 2x) = 15 x=

jigglypuff314 (jigglypuff314):

distributive property is a(b - c) = ab - ac so 4(3 - 2x) = 12 - 8x so for 4(3 - 2x) = 15 distribute 12 - 8x = 15 then subtract 12 from both sides -12 -12 -8x = 3 then divide both sides by Negative 8 /-8 /-8 so x = three divided by negative eight